Adnan scores 45% of the marks in an exam and passess by 5 marks.Taimur scores 48% of the marks in the same exam and passes by 32 marks.What are the passing marks in the exam?
Answer by robertb(5830)

Let t = maximum marks available, and p = passing marks.
Then from the first sentence, we get
0.45t = p + 5. (1)
From the 2nd sentence we obtain
0.48t = p + 32. (2)
subtracting (1) from (2), we get
0.03t = 27
t=27/0.03 = 900, the maximum marks.
Hence, p = 0.45*900 - 5 = 400, by using (1)
Therefore the passing mark is 400.